OpenStudy (aaronandyson):

(cosecA-sinA)(secA-cosA)(tanA+cotA) = 2

OpenStudy (aaronandyson):

@triciaal

OpenStudy (welshfella):

you want to prove this right?

OpenStudy (aaronandyson):

LHS = RHS

OpenStudy (welshfella):

convert to sin's and cos = ( 1/s - s)(1/c - c)( s/c + c/s) = 1 -s^2 1 - c^2 s^2 + c^2 ------ * ------- * --------- s c cs now c^2 + s^2 = 1 and 1 - s^2 = c^2 and 1 - c^2 = s^2 so we have c^2/s * s^2/c * 1/ cs = c^2 s^2 / c^2 s^2 = 1

OpenStudy (welshfella):

I make it 1 either i've made a mistake or question is wrong...

OpenStudy (aaronandyson):

Yes,It's 1 and not 2

OpenStudy (welshfella):

i expet you might now this by i used the identities sec x = 1/cos x , cot x = 1/tan x and cosex x = 1 / sin x

OpenStudy (triciaal):